# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v1

@@ 5 Classification Error of the Convergent Direction
In this section, we discuss the classification performance of the convergent direction $\hat{W}$ of (2) whose characterization given by Theorem 4.1. We establish a classification error bound by obtaining a closed form expression of $\boldsymbol{\hat{w}}=\mathrm{vec}(\hat{W})$, which is $\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$, where $\boldsymbol{y}=(y_{1},\ldots,y_{n})^{\top}\in\mathbb{R}^{n}$.
@@
$\tilde{E}_{1}(\tilde{\varepsilon}_{1},R)$ :=$:=\left\{\|ZZ^{\top}-RI_{n}\|\leq\tfrac{\tilde{\varepsilon}_{1}}{3}R\right\},$ (7)
$\tilde{E}_{2}(\tilde{\varepsilon}_{2},R)$ :=$:=\left\{\|Z\boldsymbol{\mu}\|\leq\tfrac{\tilde{\varepsilon}_{2}}{3}R\right\},$ (8)
$\tilde{E}_{3}(\tilde{\varepsilon}_{3})$ :=$:=\left\{|n_{+}-\tfrac{n}{2}|\leq\tilde{\varepsilon}_{3}\tfrac{n}{2}\right\}.$ (9)
@@
Theorem 5.1.
Suppose event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with $\tilde{\varepsilon}\leq\tfrac{q_{\gamma}}{2}$, $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$, and further assume one of the following condition holds for a sufficiently large constant $C=C(\gamma,q_{+})$:
@@
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq CR$, $C\tilde{\varepsilon}\leq\frac{R}{n^{\frac{3}{2}}\|\boldsymbol{\mu}\|^{2}}$, and $C\tilde{\varepsilon}_{2}\leq n^{-\frac{1}{2}}$.
Then, we have
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-c\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\boldsymbol{z}\|_{\psi_{2}}^{2}\{n\|\boldsymbol{\mu}\|^{2}+R\}}\right),$
where $c$ is a constant which depends on $\gamma$ and $q_{+}$.
@@
The theorem statement here is a concise versions for the purpose of presentation. Detailed version of the theorem with the explicit form of constant $C$ and the proof are given in Theorem B.11, Appendix B.
Theorem 5.1 shows that there is a phase transition between the weak signal regime, i.e. $n\|\boldsymbol{\mu}\|^{2}\lesssim R$ and the strong signal regime, i.e. $n\|\boldsymbol{\mu}\|^{2}\gtrsim R$. The weak signal regime corresponds to the nearly orthogonal data regime studied in prior works under specific distributional settings. Specifically, Frei et al. (2022) assumed $\boldsymbol{z}$ is sub-Gaussian and strongly log-concave and Xu and Gu (2023) assumed $\boldsymbol{z}$ has a bounded log-Sobolev constant, which also is stronger than assuming $\boldsymbol{z}$ is sub-Gaussian. Furthermore, both works only allow for isotropic $\boldsymbol{z}$. Our result not only extends the prior works to the strong signal regime but also establishes sufficient deterministic conditions on the data to guarantee benign overfitting, which allows us to apply the result to a wider distributional setting.
A similar phase transition in the classification error bound was previously established for binary linear classification in Cao et al. (2021); Wang and Thrampoulidis (2022); Hashimoto et al. (2025).
@@
Lastly, we note that the condition $q_{\gamma}>\gamma$ in (iii) is necessary to ensure the equivalence of $\boldsymbol{\hat{w}}$ and the minimum norm least square estimator. In fact, we can prove that the equivalence fails if $q_{\gamma}<\gamma$, which can be viewed as another form of the phase transition in the behavior of the convergent direction. Additional details are provided in Appendix B.
6 Benign Overfitting on sub-Gaussian Mixture
We here present applications of Theorem 4.4, 4.9, and 5.1 to anisotropic sub-Gaussian mixtures (sG). Our main results are given separately for the strong signal and weak signal regimes. For the strong signal regime, we can show
Theorem 6.1 (Strong Signal Regime).
Consider model (sG). Let $\delta\in(0,\tfrac{1}{4})$. Suppose $q_{\gamma}>\gamma$, $n\geq C$, $\alpha\leq\frac{1}{2.4n\|\boldsymbol{\mu}\|^{2}}$, $\sigma\leq 0.2\gamma\alpha\sqrt{\frac{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)}{m}}$,
$\|\boldsymbol{\mu}\|^{2}$ $\geq C\max\Big{\{}\frac{{\rm tr}(\Sigma)}{n},\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|,\sqrt{n}\|\Sigma\|_{F},n\|\Sigma\|\Big{\}}$
for a sufficiently large constant $C$ which depends only on $\delta,\gamma,q_{+}$. Then, we have with probably at least $1-3\delta$: the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, converges in direction, and the convergent direction $\hat{W}$ attains
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-\frac{c}{L^{2}}\frac{\|\boldsymbol{\mu}\|^{2}}{\|\Sigma\|}\right),$
where $c$ is a constant which depends only on $\gamma$ and $q_{+}$.
Theorem 6.1 establishes the test error bound for (sG) in the strong signal regime, which was not studied in the existing literature. We see that benign overfitting occurs if $\|\boldsymbol{\mu}\|^{2}\gg\|\Sigma\|$ is satisfied in addition to the assumptions of the theorem.
For the weak signal regime, we can show
Theorem 6.2 (Weak Signal Regime).
Consider model (sG). Let $\delta\in(0,\tfrac{1}{4})$. Suppose $n\geq C$, $\alpha\leq\frac{1}{8{\rm tr}(\Sigma)}$, $\sigma\leq 0.1\gamma\alpha\sqrt{\frac{{\rm tr}(\Sigma)}{m}}$
@@ 6 Benign Overfitting on sub-Gaussian Mixture
for a sufficiently large constant $C$ which depends only on $\delta,\gamma,q_{+}$. Then, we have with probably at least $1-3\delta$, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, converges in direction, and the convergent direction $\hat{W}$ attains
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-\frac{c}{L^{2}}\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\Sigma\|{\rm tr}(\Sigma)}\right),$
@@
Theorem 6.2 establishes the test error bound for (sG) in the weak signal regime, which corresponds to the prior works (Frei et al. (2022); Xu and Gu (2023)), which were limited to isotropic mixtures. We see that benign overfitting occurs if $n\|\boldsymbol{\mu}\|^{4}\gg\|\Sigma\|{\rm tr}(\Sigma)$.
As a corollary of these results, we provide the sufficient conditions of benign overfitting when $\Sigma=I_{p}$:
Corollary 6.3.
Consider model (sG) with $\Sigma=I_{p}$. Then, with high probability, the gradient descent iterate (2) converges in direction, and benign overfitting occurs with the convergent direction $\hat{W}$ if either one of the following holds:
@@
    $\|\boldsymbol{\mu}\|\gtrsim\sqrt{n}$ and $n^{\frac{5}{4}}\|\boldsymbol{\mu}\|^{\frac{3}{2}}\lesssim p\lesssim\min\{\tfrac{\|\boldsymbol{\mu}\|^{4}}{n},n\|\boldsymbol{\mu}\|^{2}\}$,
@@
We note that the sufficient condition (iii) in the corollary is directly comparable to the results studied in the prior works (Frei et al. (2022); Xu and Gu (2023)), while conditions (i) and (ii) were not identified in the existing literature. Comparing the sufficient condition (iii) with that of Xu and Gu (2023), we do not require $\boldsymbol{z}$ to have a bounded log-Sobolev constant, and we have a better upper bound on $p$ (Xu and Gu (2023) requires $p\log(mp)\lesssim n\|\boldsymbol{\mu}\|^{4}$). We also do not require $m\gtrsim\log n$, which was necessary in their work to ensure $\min\{q_{+},q_{-}\}\geq 1/3$ after random initialization of $a_{j}$’s. Although we do not consider random initialization of $a_{j}$’s, our work on the weak signal regime can incorporate the random initialization of the form $a_{j}\sim\mathrm{Unif}\left\{\pm\tfrac{1}{\sqrt{m}}\right\}$ without assuming a lower bound on $m$. We note that our lower bound on $p$ is slightly stronger than theirs which requires $p\gtrsim\max\{n\|\boldsymbol{\mu}\|^{2},n^{2}\log n\}$.
The theorem statements here are a concise versions for the purpose of presentation. Detailed versions of these theorems and the proofs are given in Appendix C.
@@ 7 Proof Sketch
Proof sketch of Theorem 4.1.
@@
where $\boldsymbol{w}^{(t)}=\mathrm{vec}(W^{(t)})$. Therefore, if all the neurons stay activated after $t\geq T$, gradient descent iterate (2) becomes that of linear classifier, and the directional convergence follows from Theorem 3, Soudry et al. (2018).
Proof sketch of Theorem 4.4 and 4.9.
@@
So, the only negativity can be controlled by $2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$. Therefore, under the condition $2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ is small enough (Case 1 or Assumption 4.5 in Case 2), the update (10) should stay positive. Then, it should be enough to make all the neurons activated at step $t=1$, which can be done by taking initialization $\sigma$ sufficiently smaller than the step size $\alpha$ (Assumption 4.2 and 4.6).
@@
To apply Lemma B.1, it is necessary to understand the behavior of $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}$. Letting $X=(\boldsymbol{x}_{1}^{\top},\boldsymbol{x}_{2}^{\top},\ldots,\boldsymbol{x}_{n}^{\top})^{\top}\in R^{n\times p}$ and recalling $\boldsymbol{\tilde{x}}_{ij}=a_{j}\zeta(y_{i}a_{j})\boldsymbol{x}_{i}$, we can write
@@
The analysis of $\boldsymbol{y}(\tilde{X}\tilde{X}^{\top})^{-1}$ will be done through $XX^{\top}=\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+\boldsymbol{y}(Z\boldsymbol{\mu})^{\top}+Z\boldsymbol{\mu}\boldsymbol{y}^{\top}+ZZ^{\top}\approx\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+RI_{n}$ under sufficient near orthogonality of $\boldsymbol{z}_{i}$’s to each other and to $\boldsymbol{\mu}$.
The rest of the argument is split into the strong signal regime and the weak signal regime. The latter is easier since we have $XX^{\top}\approx RI_{n}$, which further implies $\tilde{X}\tilde{X}^{\top}\approx R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$, which is a diagonal matrix. The strong signal case is more difficult to analyze, but this can be done by a careful analysis of the eigenvectors of $\|\boldsymbol{\mu}\|^{2}(q_{+}B_{+}\boldsymbol{y}\boldsymbol{y}^{\top}B_{+}+q_{-}B_{-}\boldsymbol{y}\boldsymbol{y}^{\top}B_{-})+R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$.

### v2

@@ 5 Classification Error of the Convergent Direction
In this section, we discuss the classification performance of the convergent direction $\hat{W}$.
@@
$\tilde{E}_{1}(\tilde{\varepsilon}_{1},R)$ $:=\left\{\|ZZ^{\top}-RI_{n}\|\leq\tfrac{\tilde{\varepsilon}_{1}}{3}R\right\},$ (6)
$\tilde{E}_{2}(\tilde{\varepsilon}_{2},R)$ $:=\left\{\|Z\boldsymbol{\mu}\|\leq\tfrac{\tilde{\varepsilon}_{2}}{3}R\right\},$ (7)
$\tilde{E}_{3}(\tilde{\varepsilon}_{3})$ $:=\left\{|n_{+}-\tfrac{n}{2}|\leq\tilde{\varepsilon}_{3}\tfrac{n}{2}\right\}.$ (8)
@@
Theorem 5.1 (Classification Error Bounds).
Suppose event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with $\tilde{\varepsilon}\leq\tfrac{q_{\gamma}}{2}$, $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$, and further assume one of the following condition holds for a constant $C=C(\gamma,q_{+})$:
@@
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq CR$, $C\tilde{\varepsilon}\leq\frac{R}{n^{\frac{3}{2}}\|\boldsymbol{\mu}\|^{2}}$.
Then, there exists constants $c,N$ which depend on $\gamma,q_{+}$ such that $n\geq N$ implies
[…]
If $\boldsymbol{z}\sim\mathcal{N}(0,\Sigma)$, then we have instead
$\kappa\left\{c^{-1}\beta_{min}^{-\frac{1}{2}}\left(\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}\right)^{\frac{1}{2}}\right\}\leq\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\kappa\left\{c\beta_{max}^{-\frac{1}{2}}\left(\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}\right)^{\frac{1}{2}}\right\},$
where $\kappa(t)=\mathbb{P}(|\xi_{1}|\geq t)$ and $\beta_{min}$ is the smallest eigenvalue of $\Sigma$ and $\beta_{max}=\|\Sigma\|$.
@@
This is a concise version for the purpose of presentation. A detailed version of the theorem and the proof are given in Theorem C.13, Appendix C.
Theorem 5.1 shows that there is a phase transition between the weak signal regime, i.e. $n\|\boldsymbol{\mu}\|^{2}\lesssim R$ and the strong signal regime, i.e. $n\|\boldsymbol{\mu}\|^{2}\gtrsim R$. The weak signal regime corresponds to the nearly orthogonal data regime studied in prior work under specific distributional settings (Frei et al. (2022); Xu and Gu (2023)). The lower bound in the case of Gaussian mixture implies that the phase transition is indeed a feature of the model. Previously, this type of phase transition was not shown even with gradient flow. Moreover, it was only known to occur for binary linear classification (Cao et al. (2021); Wang and Thrampoulidis (2022); Hashimoto et al. (2025)).
Moreover, the assumptions of Theorem 5.1 are given in a deterministic manner. We confirm in the next section that these assumptions are satisfied with high probability under model (sG) and (PM), and conjecture that they can also be verified in other settings.
@@
Lastly, we note that the condition $q_{\gamma}>\gamma$ in (iii) is necessary to ensure the equivalence of $\boldsymbol{\hat{w}}$ and the minimum norm least square estimator. We can even prove that the equivalence fails if $q_{\gamma}<\gamma$, which can be viewed as another form of the phase transition in the behavior of the convergent direction. Additional details are provided in Appendix C.
6 Benign Overfitting on Mixture Models
We here demonstrate that the results of Theorem 4.8 and 5.1 hold with high probability in model (sG) and (PM) under sufficient over-parametrization. For (sG) we make the following assumptions:
Assumption 6.1.
One of the following conditions holds for a constant $C$
  1. (a)
    $n\geq C$, $\alpha\leq\frac{1}{8{\rm tr}(\Sigma)}$, $\sigma\leq 0.1\gamma\alpha\sqrt{\frac{{\rm tr}(\Sigma)}{m}}$, $\|\boldsymbol{\mu}\|^{2}\geq C\|\Sigma^{]}frac{1}{2}\boldsymbol{\mu}\|$,
@@ 6 Benign Overfitting on Mixture Models
  2. (b)
    $q_{\gamma}>\gamma$, $n\geq C$, $\alpha\leq\frac{1}{2.4n\|\boldsymbol{\mu}\|^{2}}$, $\sigma\leq 0.2\gamma\alpha\sqrt{\frac{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)}{m}}$,
    $\|\boldsymbol{\mu}\|^{2}$ $\geq C\max\Big\{\frac{{\rm tr}(\Sigma)}{n},\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|,\sqrt{n}\|\Sigma\|_{F},n\|\Sigma\|\Big\}.$
We note that condition (a) above corresponds to the weak signal regime and (b) to the strong signal regime in Theorem 5.1. We are now ready to present our main classification error bound on (sG):
Theorem 6.2.
Consider model (sG). Let $\delta\in(0,\frac{1}{4})$. Suppose Assumption 6.1 holds for a sufficiently large constant $C$ which depends only on $\delta,\gamma,q_{+}$. Then, we have with probably at least $1-3\delta$: the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, converges in direction, and the convergent direction $\hat{W}$ attains
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-\frac{c}{L^{2}}\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\Sigma\|\{n\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}}\right),$
@@
If $\boldsymbol{z}\sim\mathcal{N}(0,\Sigma)$, then we have instead
$\kappa\left\{c^{-1}\beta_{min}^{-\frac{1}{2}}\left(\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)}\right)^{\frac{1}{2}}\right\}\leq\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\kappa\left\{c\beta_{max}^{-\frac{1}{2}}\left(\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)}\right)^{\frac{1}{2}}\right\}.$
Theorem 6.2 implies that benign overfitting occurs if $n\|\boldsymbol{\mu}\|^{4}\gg\|\Sigma\|{\rm tr}(\Sigma)$ under (a) or $\|\boldsymbol{\mu}\|^{2}\gg\|\Sigma\|$ under (b) is satisfied. For Gaussian mixture model, we see that benign overfitting provably fails even though directional convergence occurs if $n\|\boldsymbol{\mu}\|^{4}\lesssim\beta_{min}{\rm tr}(\Sigma)$ under (a) or $\|\boldsymbol{\mu}\|^{2}\lesssim\beta_{min}$ under (b).
We note that Theorem 6.2 under Assumption 6.1 (a) (i) and (b) are novel regimes which were not identified in the existing literature (Frei et al. (2022); Xu and Gu (2023)). Only the result under Assumption 6.1 (a) (ii) with $\Sigma=I_{p}$ was previously studied. A more detailed comparison with the prior work is given in Appendix A.
Moreover, Theorem 4.8 and 5.1 allow us to extend the above result to a heavier tailed model (PM). This is a significant distributional relaxation from the existing results which assumed $\boldsymbol{z}$ is sub-Gaussian and strongly log-concave (Frei et al. (2022)) or $\boldsymbol{z}$ has a bounded log-Sobolev constant (Xu and Gu (2023)). For brevity, we only present a result for $\Sigma=I_{p}$:
Theorem 6.3.
Consider model (PM) with $\Sigma=I_{p}$. Then, for sufficiently large $n$, benign overfitting occurs with the convergent direction $\hat{W}$ with high probability if either one of the following holds:
@@
    $\|\boldsymbol{\mu}\|\gtrsim\sqrt{n}$ and $\max\{n\|\boldsymbol{\mu}\|^{2},n^{\frac{r+4}{2r-4}},n^{\frac{8}{r}+1}\}\lesssim p\lesssim\min\left\{\frac{\|\boldsymbol{\mu}\|^{4}}{n^{\frac{8}{r}}},\frac{\|\boldsymbol{\mu}\|^{r}}{n}\right\}$,
@@
    $\|\boldsymbol{\mu}\|\gtrsim\sqrt{n}$ and
    $\max\{n^{\frac{5}{4}}\|\boldsymbol{\mu}\|^{\frac{3}{2}},n^{\frac{8}{3r}+1}\|\boldsymbol{\mu}\|^{\frac{4}{3}},n^{\frac{3r+4}{3r-4}}\|\boldsymbol{\mu}\|^{\frac{4r}{3r-4}}\}\lesssim p\lesssim\min\left\{n\|\boldsymbol{\mu}\|^{2},\frac{\|\boldsymbol{\mu}\|^{4}}{n^{\frac{8}{r}}},\frac{\|\boldsymbol{\mu}\|^{r}}{n}\right\}$,
@@
    $\max\{n\|\boldsymbol{\mu}\|^{2},n^{\frac{r+4}{2r-4}},n^{\frac{8}{r}+2}\}\lesssim p\ll n\|\boldsymbol{\mu}\|^{4}$.
Detailed versions of the theorems and the proofs are given in Appendix D.
@@ 7 Proof Sketch
In this section, we provide the proof sketch of our main theorems. The key to establish the classification error bound in Theorem 5.1 is the precise characterization of the convergent direction in Theorem 4.8. The precise characterization allows us to obtain both upper and lower bounds of the key quantity $\frac{\|\boldsymbol{\bar{w}}\|}{\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle}$ in classification error bounds (98), which results in the novel lower bound in the case of Gaussian mixture. We also note that the conditions in Theorem 4.8 and 5.1 are given in a deterministic manner. The separation of deterministic and distributional arguments allows us to substantially relax distributional assumptions made in prior work. In fact, Theorem 6.2 and 6.3 are proved by showing that all the conditions in Theorem 4.8 and 5.1 are achieved with high probability under (sG) and (PM) with sufficient overparametrization.
Proof sketch of Theorem 4.8
The key to show directional convergence of the gradient descent is activation of neurons. We say that $j$-th neuron $\phi(\langle\cdot,\boldsymbol{w}_{j}\rangle)$ is activated if $a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle)>0$ holds for any $i\in I$, which can be equivalently written as $a_{j}y_{i}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle>0$ for any $i\in I$.
[…]
Lemma 7.1.
[…]
@@
where $\boldsymbol{w}^{(t)}=\mathrm{vec}(W^{(t)})$. Therefore, if all the neurons stay activated after $t\geq T$, gradient descent iterate (2) becomes that of linear classifier in a transformed space, and the directional convergence follows from Theorem 3, Soudry et al. (2018).
@@
So, by keeping $2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ small (Assumption 4.1 and 4.4), the update (10) should stay positive. Then, it should suffice to make all the neurons activated at step $t=1$, which can be done by taking initialization $\sigma$ sufficiently smaller than the step size $\alpha$ (Assumption 4.2 and 4.5).
@@
[…]
@@
The analysis of $\boldsymbol{y}(\tilde{X}\tilde{X}^{\top})^{-1}$ will be done through $XX^{\top}=\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+\boldsymbol{y}(Z\boldsymbol{\mu})^{\top}+Z\boldsymbol{\mu}\boldsymbol{y}^{\top}+ZZ^{\top}\approx\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+RI_{n}$ under sufficient near orthogonality of $\boldsymbol{z}_{i}$’s to each other and to $\boldsymbol{\mu}$ (events $\tilde{E}_{1},\tilde{E}_{2}$).
The rest of the argument is split into the strong signal regime and the weak signal regime. The latter is easier since we have $XX^{\top}\approx RI_{n}$, which further implies $\tilde{X}\tilde{X}^{\top}\approx R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$, which is a diagonal matrix. The strong signal case is more difficult to analyze, and relies on a careful characterization of the eigenvectors of $\|\boldsymbol{\mu}\|^{2}(q_{+}B_{+}\boldsymbol{y}\boldsymbol{y}^{\top}B_{+}+q_{-}B_{-}\boldsymbol{y}\boldsymbol{y}^{\top}B_{-})+R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$.